It is 260 km from Patna to Ranchi by air and 320 km by road. An aeroplane takes 30 minutes to go from Patna to Ranchi whereas a delux bus takes 8 hours. (a) Find the average speed of the plane. (b) Find the average speed of the bus. (c) Find the average velocity of the plane. (d) Find the average velocity of the bus.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. (a) Vave of plane (Distance / Time) = 260/0.5 = 520 km/hr. (b) Vave of bus = 320 / 8 = 40 km/hr. d) Straight path distance between plane to Ranchi is equal to the displacement of bus. ∴ Velocity = V ⃗ave = 260/8 = 32.5 km/hr.